# Boltzmann to Lindblad: Classical and Quantum Approaches to Out-of-Equilibrium Statistical Mechanics

### v1

@@ 1 Introduction
A major advance came with the formulation of the quantum master equation, describing the evolution of an open system’s density matrix [8, 9, 10, 11]. The Redfield equation, valid for weak system–environment coupling [8], found early applications in nuclear magnetic resonance but does not ensure positivity of the density matrix [12]. This limitation motivated the development of more rigorous approaches, including path-integral formulations, quantum thermodynamics, and quantum Brownian motion [13, 14, 15, 16, 17, 18, 19, 20]. In the 1970s, these efforts culminated in the Gorini–Kossakowski–Sudarshan–Lindblad (GKLS) equation [21, 22, 23, 24, 25], which provides a fully Markovian and mathematically consistent framework for quantum dynamical semigroups, refining the Redfield approximation [8], and advancing the theoretical foundations of quantum thermodynamics [26, 27, 28, 29, 30, 31, 32, 33]. This framework clarified the mechanisms of decoherence and environmental coupling [34, 35, 36] and inspired subsequent studies on quantum equipartition [37, 38, 39, 40, 41] and fluctuation theorems [42, 43].
@@
Today, open quantum system theory constitutes a cornerstone of modern physics. It underpins applications in nuclear magnetic resonance [44, 45], optical pumping [46], masers [47, 48], and lasers [49, 50], as well as in quantum computation [51, 52, 53, 54], quantum information [55, 56], and molecular electronics [57, 58]. The same framework informs quantum biology [59, 60, 61, 62], notably in photosynthetic energy transfer [63, 64, 65], and plays a central role in quantum thermodynamic devices [66], quantum batteries [67], heat engines [68], and in understanding the thermodynamic arrow of time [69, 70, 71]. Furthermore, it provides the theoretical foundation for quantum metrology and sensing [72, 73, 74], optomechanics [75, 76], nonequilibrium phenomena in nanoscale and condensed-matter systems [77, 78, 79, 80], and engineering of active quantum matter [81].
@@
In the present work, which extends the analysis of Ref. [82], we firstly construct a generalized classical Langevin equation [83], in which both friction and noise appear symmetrically in the Hamiltonian formalism. By using the Fokker–Planck methodology [84, 85], we derive a generalized Klein–Kramers equation [86, 87], expressed in terms of Poisson brackets, and show that the asymptotic solution of this equation corresponds to the classical canonical distribution. Moreover, by introducing the definitions of heat [88, 89], and entropy [90, 91, 92], along a single trajectory, we demonstrate that this approach is fully consistent with the first and second laws of thermodynamics. It is worth noting that such a formalism, based on the theory of stochastic processes [93, 94, 95, 98, 99], is now systematically applied within statistical mechanics [100, 101, 102, 103, 104, 105], including holonomic underdamped and overdamped systems [108, 109, 106, 107]. Secondly, we introduce the quantum analog of the classical theory. We build on our previous studies [110, 111, 112, 113], by means of which we showed that noise in quantum mechanics always enters multiplicatively, and demonstrated that canonical quantization – i.e., replacing Poisson brackets with commutator [114, 115] – constitutes a rigorous procedure in the classical-to-quantum transition [82]. This approach has also been discussed and/or applied in various studies [116, 117, 118, 119, 120, 121]. Accordingly, we apply canonical quantization to the Klein–Kramers equation, which is based on the symmetric inclusion of friction and noise in the Hamiltonian equations. This procedure necessitates the introduction of certain friction operators, which can be defined in two distinct ways depending on whether Hermiticity is imposed or not. By following both approaches in parallel, we thus obtain two different master equations to describe the nonequilibrium statistical mechanics of a quantum system. For both master equations, we analyze their behavior for a simple system, such as the harmonic oscillator. This analysis, based on Refs. [117, 118], shows that the resulting equations are of the Lindblad type, i.e. they are completely positive, only if both Hamiltonian equations include friction and noise terms. This fully justifies the initial choice made in reconstructing the classical statistical mechanics, where friction and noise were introduced in both Hamiltonian equations. Furthermore, we demonstrate that in the quantum case the two always-positive friction coefficients must also satisfy an additional inequality to ensure complete positivity. Remarkably, this inequality is identical for both the Hermitian and non-Hermitian operator approaches, suggesting that it could be a more general feature rather than a consequence of the specific choice of friction operators. In the following, we show how quantum thermodynamics can be developed starting from the master equation, leading to a novel formulation of the first and second laws. The first law is based on a generalized definition of heat at the quantum level, while the second law asserts the positivity of entropy production, highlighting its deep connection with the monotonicity of quantum relative entropy [21, 22]. It should be noted that multiple approaches have been developed to investigate the positivity of entropy production in the quantum context [26, 27, 28, 29, 30, 31].
@@ 2 Generalized classical stochastic thermodynamics
The symmetry of this Hamiltonian form justifies the choice made to have a form of friction in Eq. (2) proportional to $\frac{\partial V_{0}}{\partial\vec{r}_{i}}=\frac{\partial\mathcal{H}_{0}}{\partial\vec{r}_{i}}$. Further examples of generalized Hamilton–Langevin formulations are available in the existing literature [96, 97]. From a mathematical point of view, our peculiar writing of the Hamilton equations represents a stochastic differential problem with additive noise [98, 99]. We can then apply the Fokker-Planck methodology, which is briefly recalled here for an arbitrary interpretation of the stochastic calculus [93, 94, 95]. Although this distinction is strictly speaking not relevant here since the classical Hamilton equations have an additive noise, we underline that the corresponding quantum equations have multiplicative noise, as thoroughly discussed in Refs. [116, 82]. We thus consider the stochastic differential system
@@ 4.2 Relation with the quantum optical master equation
We notice that this result is in perfect agreement with what we found in the previous subsection about the conditions for the Lindblad structure of Eq. (4.1). In particular, Eq. (4.1) coincides with the QOME in Eq. (4.2) for the choice $\mu=0$ in Eq. (49). Importantly, it means that the QOME in Eq. (4.2) is always represented by a point lying on the bisector represented in Fig. 1. Thus, the results obtained for the damped harmonic oscillator in the quantum framework, described in the previous section, are fully coherent with the outcomes used in the literature based on the QOME.
@@ 6 Quantum stochastic thermodynamics
In this section, we construct the thermodynamic formulation starting from the evolution equation for the density matrix introduced in Eq. (23). For the sake of clarity and without loss of generality, we confine our discussion to the case of Hamiltonian friction operators, which are not only analytically tractable but also the most relevant from a physical standpoint. Within this framework, we establish the first and second laws of thermodynamics after introducing the notions of internal energy and entropy associated with the system.
@@ 6.2 Second law of thermodynamics
which means that the rate of entropy production can be determined through the new expression:
$\frac{d\mathcal{S}_{p}}{dt}=\left.-k_{B}\frac{d}{d\tau}\mathcal{S}\left(e^{\mathcal{L}\tau}\varrho|\varrho_{eq}\right)\right|_{\tau=0}.$ (98)
The latter can be further elaborated as follows:
@@
It should be noted that the non-negativity of the quantum entropic production is discussed in several works, with different approaches and methodologies [26, 27, 28, 29, 30, 31].
@@ 8 Conclusions
A natural perspective for future research is to establish that the symmetry of friction and noise in the Hamilton equations, for arbitrary potential energy landscapes, systematically leads—through canonical quantization—to master equations of Lindblad type, and therefore to completely positive quantum evolutions. This question can be approached by analyzing additional simple models for which the energy basis allows fully analytic expressions for all operators, such as the infinite potential well, and then extending the analysis to general potential energies. Depending on the complexity of the potential, this program may be pursued analytically or numerically.

### v2

@@ 1 Introduction
A major advance came with the formulation of the quantum master equation, describing the evolution of an open system’s density matrix [8, 9, 10, 11]. The Redfield equation, valid for weak system–environment coupling [8], found early applications in nuclear magnetic resonance but does not ensure positivity of the density matrix [12]. This limitation motivated the development of more rigorous approaches, including path-integral formulations, quantum thermodynamics, and quantum Brownian motion [13, 14, 15, 16, 17, 18, 19, 20]. In the 1970s, these efforts culminated in the Gorini–Kossakowski–Sudarshan–Lindblad (GKLS) equation [21, 22, 23, 24, 25], which provides a fully Markovian and mathematically consistent framework for quantum dynamical semigroups, refining the Redfield approximation [8], and advancing the theoretical foundations of quantum thermodynamics [26, 27, 28, 29, 30, 31, 32, 33, 34]. This framework clarified the mechanisms of decoherence and environmental coupling [35, 36, 37] and inspired subsequent studies on quantum equipartition [38, 39, 40, 41, 42] and fluctuation theorems [43, 44].
@@
Today, open quantum system theory constitutes a cornerstone of modern physics. It underpins applications in nuclear magnetic resonance [45, 46], optical pumping [47], masers [48, 49], and lasers [50, 51], as well as in quantum computation [52, 53, 54, 55], quantum information [56, 57], and molecular electronics [58, 59]. The same framework informs quantum biology [60, 61, 62, 63], notably in photosynthetic energy transfer [64, 65, 66], and plays a central role in quantum thermodynamic devices [67], quantum batteries [68], heat engines [69], and in understanding the thermodynamic arrow of time [70, 71, 72]. Furthermore, it provides the theoretical foundation for quantum metrology and sensing [73, 74, 75], optomechanics [76, 77], nonequilibrium phenomena in nanoscale and condensed-matter systems [78, 79, 80, 81], and engineering of active quantum matter [82]. In many realistic physical situations, the interaction between a system and its environment gives rise to memory effects, making non-Markovian descriptions of open quantum dynamics both necessary and physically meaningful [83, 84, 85, 86].
@@
In the present work, which extends the analysis of Ref. [87], we firstly construct a generalized classical Langevin equation [88], in which both friction and noise appear symmetrically in the Hamiltonian formalism. By using the Fokker–Planck methodology [89, 90], we derive a generalized Klein–Kramers equation [91, 92], expressed in terms of Poisson brackets, and show that the asymptotic solution of this equation corresponds to the classical canonical distribution. Moreover, by introducing the definitions of heat [93, 94], and entropy [95, 96, 97], along a single trajectory, we demonstrate that this approach is fully consistent with the first and second laws of thermodynamics. It is worth noting that such a formalism, based on the theory of stochastic processes [98, 99, 100, 101, 102], is now systematically applied within statistical mechanics [103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113], including holonomic underdamped and overdamped systems [116, 117, 114, 115]. Secondly, we introduce the quantum analog of the classical theory. We build on our previous studies [118, 119, 120, 121], by means of which we showed that noise in quantum mechanics always enters multiplicatively, and demonstrated that canonical quantization – i.e., replacing Poisson brackets with commutator [122, 123] – constitutes a rigorous procedure in the classical-to-quantum transition [87]. This approach has also been discussed and/or applied in various studies [124, 125, 126, 127, 128, 129]. Accordingly, we apply canonical quantization to the Klein–Kramers equation, which is based on the symmetric inclusion of friction and noise in the Hamiltonian equations. This procedure necessitates the introduction of certain friction operators, which can be defined in two distinct ways depending on whether Hermiticity is imposed or not. By following both approaches in parallel, we thus obtain two different master equations to describe the nonequilibrium statistical mechanics of a quantum system. For both master equations, we analyze their behavior for a simple system, such as the harmonic oscillator. This analysis, based on Refs. [125, 126], shows that the resulting equations are of the Lindblad type, i.e. they are completely positive, only if both Hamiltonian equations include friction and noise terms. This fully justifies the initial choice made in reconstructing the classical statistical mechanics, where friction and noise were introduced in both Hamiltonian equations. Furthermore, we demonstrate that in the quantum case the two always-positive friction coefficients must also satisfy an additional inequality to ensure complete positivity. Remarkably, this inequality is identical for both the Hermitian and non-Hermitian operator approaches, suggesting that it could be a more general feature rather than a consequence of the specific choice of friction operators. In the following, we show how quantum thermodynamics can be developed starting from the master equation, leading to a novel formulation of the first and second laws. The first law is based on a generalized definition of heat at the quantum level, while the second law asserts the positivity of entropy production, highlighting its deep connection with the monotonicity of quantum relative entropy [21, 22]. It should be noted that multiple approaches have been developed to investigate the positivity of entropy production in the quantum context [26, 27, 28, 29, 30, 31, 32].
@@ 2 Generalized classical stochastic thermodynamics
The symmetry of this Hamiltonian form justifies the choice made to have a form of friction in Eq. (2) proportional to $\frac{\partial V_{0}}{\partial\vec{r}_{i}}=\frac{\partial\mathcal{H}_{0}}{\partial\vec{r}_{i}}$. Further examples of generalized Hamilton–Langevin formulations are available in the existing literature [109, 110, 111, 112, 113]. From a mathematical point of view, our peculiar writing of the Hamilton equations represents a stochastic differential problem with additive noise [101, 102]. We can then apply the Fokker-Planck methodology, which is briefly recalled here for an arbitrary interpretation of the stochastic calculus [98, 99, 100]. Although this distinction is strictly speaking not relevant here since the classical Hamilton equations have an additive noise, we underline that the corresponding quantum equations have multiplicative noise, as thoroughly discussed in Refs. [124, 87]. We thus consider the stochastic differential system
@@ 3 Density matrix evolution in quantum mechanics
Here, for the sake of clarity and simplicity, all properties are presented assuming a finite ensemble of $M$ pure states. We emphasize, however, that the formalism is fully general and can be extended to more general statistical mixtures. In particular, the derivations remain valid for countably infinite ensembles, where the sum over the states extends to infinity, provided the probabilities converge so that the density matrix remains trace-class. Furthermore, the results also hold for continuous ensembles, in which the discrete sum is replaced by an integral over a continuous set of states. In both cases, the key properties of the density matrix — Hermiticity, positivity, and unit trace — are preserved, ensuring that all conclusions drawn for finite ensembles carry over straightforwardly to these more general settings.
@@ 4.2 Relation with the quantum optical master equation
We notice that this result is in perfect agreement with what we found in the previous subsection about the conditions for the Lindblad structure of Eq. (4.1). In particular, Eq. (4.1) coincides with the QOME in Eq. (4.2) for the choice $\mu=0$ in Eq. (49). Importantly, it means that the QOME in Eq. (4.2) is always represented by a point lying on the bisector represented in Fig. 1. Thus, the results obtained for the damped harmonic oscillator in the quantum framework, described in the previous section, are fully coherent with the outcomes used in the literature based on the QOME. We also note that the cavity model introduced in Ref. [86] exhibits a clear analogy with our approach based on symmetric friction and noise in the Langevin equations, and is indeed consistent with the optical master equation.
@@ 6 Quantum stochastic thermodynamics
In this section, we construct the thermodynamic formulation starting from the evolution equation for the density matrix introduced in Eq. (23). For the sake of clarity and without loss of generality, we confine our discussion to the case of Hamiltonian friction operators, which are not only analytically tractable but also the most relevant from a physical standpoint. Within this framework, we establish the first and second laws of thermodynamics after introducing the notions of internal energy and entropy associated with the system. The same line of analysis could be extended to the second case with non-Hermitian friction operators; nevertheless, we refrain from presenting it here for conciseness, as it yields results essentially analogous to those described below.
@@ 6.2 Second law of thermodynamics
which means that the rate of entropy production can be determined through the expression:
$\frac{d\mathcal{S}_{p}}{dt}=\left.-k_{B}\frac{d}{d\tau}\mathcal{S}\left(e^{\mathcal{L}\tau}\varrho|\varrho_{eq}\right)\right|_{\tau=0},$ (98)
which has been first proved in Ref.[26]. The latter can be further elaborated as follows:
@@
We note that the non-negativity of quantum entropy production has been established in the literature through a variety of approaches and methodologies, which nonetheless lead to conclusions consistent with the result obtained here [26, 27, 28, 29, 30, 31, 32].
@@ 8 Conclusions
A natural perspective for future research is to establish that the symmetry of friction and noise in the Hamilton equations, for arbitrary potential energy landscapes, systematically leads—through canonical quantization—to master equations of Lindblad type, and therefore to completely positive quantum evolutions. This question can be approached by analyzing additional simple models for which the energy basis allows fully analytic expressions for all operators, such as the infinite potential well, and then extending the analysis to general potential energies. Depending on the complexity of the potential, this program may be pursued analytically or numerically. Another promising line of investigation concerns the generalization of the present models to non-Markovian dynamics, in which memory effects become relevant.